# Spectral Criteria for the Asymptotics of Local Functionals of Gaussian Fields and Their Application to Nodal Volumes

### v2

@@ 5.2 Proofs of the variance of the nodal measure
According to Lemma 4.4, the spectral density $\Sigma$ has expression
$\Sigma(\xi)=\psi(\xi)\,\otimes\left[(1,2i\pi\xi)\otimes(1,-2i\pi\xi)\right].$
@@
It follows from Lemma 4.5 that the second chaotic component $f_{2}$ has the expression, in an orthogonal basis on $W$
$f_{2}=\mathrm{Id}\otimes\begin{pmatrix}[c|c]-\alpha&0\\ \hline\cr 0&\frac{\alpha}{d}\,\mathrm{Id}\end{pmatrix}.$
Using the definition of the trace, or equivalently the mixed product property, and the second part of Lemma 3.2 applied to the rank one matrix $(1,2i\pi\xi)\otimes(1,-2i\pi\xi)$, one gets
@@
If $Y$ is not a random wave then this quantity is positive on a set of positive $\nu$-measure, and the conclusion follows from the proof Lemma 3.11. Conversely, if $Y$ is a regular random wave then this quantity cancels for all $\xi$ in the support of $\nu$ and the conclusion of Theorem 1.8 follows from the proof of Lemma 3.12, and in particular the second point.
@@
If $U=\mathbb{R}$, the result follows directly from the previous hypothesis. If $k>1$ one can assume that, up to a change of coordinates on $V$, $\nabla\langle Y,u\rangle$ is a standard Gaussian. Let $\psi_{u}$ be the spectral density of $\langle Y,u\rangle$. The spectral density $\Sigma$ is block diagonal with block $\Sigma_{u}$ and $\Sigma_{u^{\perp}}$, where
$\Sigma_{u}(\xi)=\psi_{u}(\xi)\left[(1,2i\pi\xi)\otimes(1,-2i\pi\xi)\right].$
By independence, it follows from Lemma 4.5 that the second chaotic component is block diagonal with blocks $f_{2,u}$ and $f_{2,u^{\perp}}$, where
$f_{2,u}=\begin{pmatrix}[c|c]-\alpha&0\\ \hline\cr 0&\alpha M\end{pmatrix}$
and $M$ is a matrix (it is actually a diagonal matrix with trace $d$). As in the proof for hypothesis $(H1)$, one has
$\operatorname{Tr}\left(f_{2}\Sigma(\xi)f_{2}\Sigma(\xi)\right)\geq\left(-\alpha+4\pi^{2}\frac{\alpha}{d}\,\xi^{T}M\xi\right)^{2}|\psi_{u}(\xi)|^{2},$
@@
Let $F$ a continuous Gaussian field from $V$ to $\mathbb{R}$ with covariance function $q$ and spectral measure $\omega$, such that $Y=\nabla F$. The field $Y$ is taking values in $V$ and one can assume by a proper rescaling that $r(0)=\operatorname{Hess}q(0)=\mathrm{Id}$. Then, refining expression (5.2.2), one has
$\Sigma(\xi)=\omega(\xi)\left[(2i\pi\xi)\otimes(-2i\pi\xi)\right]\otimes\left[(1,2i\pi\xi)\otimes(1,-2i\pi\xi)\right].$
As for the second chaotic component $f_{2}$, it can be written in an orthonormal basis as
$f_{2}=\begin{pmatrix}[c|c]-\alpha\,\mathrm{Id}&0\\ \hline\cr 0&H\end{pmatrix},$
for some matrix $M$. As pointed out in Remark 4.2 we have to work with the rescaled field $\tilde{X}=\Pi X$. Because $Y(0)$ is already a standard Gaussian vector, independent from $\nabla Y(0)$, the mapping $\Pi$ can be chosen so that it only acts on the part $V\otimes V$ of $W$. The mapping $\tilde{\Sigma}$ has rank one, and it follows again from Lemma 3.2 that
$\operatorname{Tr}\left(f_{2}\tilde{\Sigma}(\xi)f_{2}\tilde{\Sigma}(\xi)\right)=\omega(\xi)^{2}\left(-4\pi^{2}\xi^{2}\alpha+(\Pi(2i\pi\xi)\otimes(2i\pi\xi))^{T}H\Pi(-2i\pi\xi)\otimes(-2i\pi\xi)\right)^{2}.$
The conclusion follows in a similar fashion as in the previous case.
@@
For $u$-level with $u\neq 0$, one can prove by similar arguments that there is no second chaos cancellation for the isotropic random waves model. Indeed, Equation (5.2.2) becomes, thanks to Lemma 4.5,
@@
Let $\mu$ be the spectral measure of the process $(Y,\nabla Y)$ taking values in the space of Hermitian forms on $W=U\otimes(\mathbb{R}\times V)$. We want to apply Theorem 1.4 with $q=2$, i.e. to show that $\operatorname{Im}(\mu^{*2})\notin C(f_{4})$ and conclude that the variance converges to a limit constant. The case $d=2$ is a bit different because of the logarithmic singularity in Lemma 4.8, but the proof follows the same line as in the case $d\geq 3$ with minor adaptations that will be detailed.
From Lemma 4.4, the spectral measure $\mu$ has a density w.r.t the spherical measure $\sigma$ given by
$\mathrm{d}\mu(\xi)=\mathrm{Id}\otimes\left[(1,2i\pi\xi)\otimes(1,-2i\pi\xi)\right]\mathrm{d}\sigma(\xi),$
Thanks to Lemma 4.6, the measure $\mu*\mu$ has a density $\Sigma$ w.r.t the measure $\sigma*\sigma$, which is bounded and polar-continuous. In light of 1.4 with $q=2$, we want to show that $\operatorname{Im}(\Sigma)\notin C(f_{4})$. Since $\Sigma$ is polar-continuous, it is enough to show the following lemma, whose proof is postponed to the end of the present proof. {mdframed}[backgroundcolor=white]
Lemma 5.2.
For $\theta\in\mathcal{S}^{d-1}$, one has
$\operatorname{Im}(\Sigma(0,\theta))\not\subset C(f_{4}).$
We distinguish the cases $d=2$ and $d\geq 3$.
Case $d\geq 3$:
In that case, $\Omega\in L^{4}$ and it follows from Theorem 1.5 that the variance of $Z_{\lambda}^{(4+)}(\phi)$ converges to a finite constant, which is positive thanks to Lemma 5.2 and Lemma 1.4, and satisfies a CLT. Since the variance of $Z_{\lambda}^{(2)}(\phi)$ converges to zero, then $Z_{\lambda}^{(2)}(\phi)$ converges in probability to zero and the CLT for $Z_{\lambda}(\phi)$ follows from Slutsky’s Theorem.
Case $d=2$:
The variance of the fourth chaos has the expression
$\operatorname{Var}(Z_{\lambda}^{(4)}(\phi))$ $=\frac{1}{24}\int_{\mathbb{R}^{d}}\left(|\hat{\phi}|^{2}\right)_{\lambda}(\xi)\,\mathrm{d}(\mu*\mu)^{*2}(f_{4})(\xi)$
$=\frac{1}{24}\int_{\mathbb{R}^{d}}\left(|\hat{\phi}|^{2}\right)_{\lambda}(\xi)\,\mathrm{d}\left[\Sigma(\sigma^{*2})*\Sigma(\sigma^{*2})\right](f_{4})(\xi)$
@@
The variance asymptotics is explicit thanks to Lemma 4.8. By integration of equivalences, it is given for some positive constant $C,C^{\prime}$ by
@@
$\lim_{\lambda\rightarrow+\infty}\frac{\operatorname{Var}(Z_{\lambda}^{(4)}(\phi))}{\log(\lambda)}=C\int_{\mathcal{S}^{1}}\Sigma(0,\theta)^{\otimes 2}(f_{4})\mathrm{d}\theta+C^{\prime}\int_{\|\xi\|=\frac{\sqrt{d}}{\pi}}\Sigma(\xi)^{\otimes 2}(f_{4})\mathrm{d}x.$
@@
Note that both terms are non-negative. From Lemma 5.2 and Lemma 3.3, the integrand in the first term is positive and the conclusion follows. Note that one can completely explicit the Hermitian form $\Sigma$ (see for instance the following proof of Lemma 5.2), and the fourth chaotic projection $f_{4}$ (by similar computation as in Lemma 4.5) to explicit the limiting constant. The CLT for $Z_{\lambda}^{(4)}(\phi)$ follows directly from the fourth Moment theorem, since $\Omega\in L^{6}$ implies that the contractions tends to zero. The variance of $Z_{\lambda}^{(6+)}(\phi)$ converges to a constant, and so does the variance of $Z_{\lambda}(\phi)-Z_{\lambda}^{(4)}(\phi)$. The CLT for $Z_{\lambda}(\phi)$ follows from Slutsky’s Theorem.
Proof of Lemma 5.2 :
Proof.
Assume first that $U=\mathbb{R}$, i.e. $k=1$. The mapping $\Sigma$ acts on the space of Hermitian forms on the complexification of the space $W=\mathbb{R}\times V$. Let $e_{0},e_{1},\ldots,e_{d}$ be a basis of ×$\mathbb{R}\times V$.The choice of $\theta$ is unimportant by isotropy of the problem and one can assume that $\theta=e_{2}$. We write the mapping $\Sigma(0,e_{2})$ in the induced basis on $V$. For $0\leq i\leq j\leq d$ and $0\leq k\leq l\leq d$, we define
$\Sigma_{(ij)(kl)}=\Sigma(0,e_{2})(\operatorname{Sym}(e_{i}\otimes e_{j}),\operatorname{Sym}(e_{k}\otimes e_{l})),$
All these quantities can be computed thanks to the second part of Lemma 4.6 : let $\rho_{0}=1$ and for $1\leq j\leq$, let $\rho_{j}:x\mapsto 2i\pi\langle x,e_{j}\rangle$. If we define $B_{1}$ the $d-2$-sphere of radius $1$, normal vector $e_{2}$ and uniform probability measure $\sigma_{2}$, then up to a normalization constant $C$ one has
$\Sigma_{(ij)(kl)}=C\int_{B_{1}}\rho_{i}(x)\rho_{j}(x)\overline{\rho_{k}(x)}\overline{\rho_{l}(x)}\mathrm{d}\;\sigma_{2}(x).$
One can explicitly compute all the coefficient of $\Sigma$ but it won’t be necessary : all we need is to observe that $\Sigma_{(01)(01)}>0$, and for $(ij)\neq(01)$, one has $\Sigma_{(01)(ij)}=0$ by symmetry. It implies that the unit vector $\operatorname{Sym}(e_{0}\otimes e_{1})$ is in the image of $\Sigma(0,\theta)$. Then
$(f_{4})(\operatorname{Sym}(e_{0}\otimes e_{1}),\operatorname{Sym}(e_{0}\otimes e_{1}))$ $=f_{4}(e_{0}\otimes e_{0}\otimes e_{1}\otimes e_{1})$
$=\mathbb{E}[f(X(0))(Y(0)^{2}-1)((\partial_{1}Y(0))^{2}-1)]$
$=\zeta$
where the last equality follows from Lemma 4.5. Since $\zeta\neq 0$, it implies that $\operatorname{Sym}(e_{0}\otimes e_{1})\notin C(f_{4})$ and thus that $\operatorname{Im}(\Sigma(0,\theta))\not\subset C(f_{4})$, which proves the lemma in the case $k=1$.
For general $U$, the mapping $\Sigma$ acts on the space of Hermitian forms on the complexification of the space $W=U\otimes(\mathbb{R}\times V)$. Let $u_{1},\ldots,u_{k}$ be a basis of $U$. Notice first that the action of mapping $\Sigma$ can be decomposed as the sum of its actions on each space $W_{i}=\{u_{i}\}\otimes(\mathbb{R}\times V)$, by independence of the coordinates. It then suffices to apply the reasoning above reasoning for $k=1$, to the restriction of $\Sigma$ to any of these spaces to show that the unity vector $\operatorname{Sym}((u_{i}\otimes e_{0})\otimes(u_{i}\otimes e_{1}))$ is in the image of $\Sigma(0,e_{2})$, but not in the isotropic cone $C(f_{4})$, and the conclusion follows. ∎

### v3

@@ 5.2 Proofs of the variance of the nodal measure
5.2.2 Proof of Theorem 1.6
Using the independence between $Y(0)$ and $\nabla Y(0)$ and the symmetry of the Jacobian map, the first chaotic projection for the nodal volume has the expression
$f_{1}=\alpha_{u}(u,0).$
Using Corollary 1.3 with $q=1$ we deduce that
$\lim_{\lambda\rightarrow+\infty}\operatorname{Var}(Z_{\lambda}^{u,(1)}(\phi))=\alpha_{u}\Sigma(0)(u,u).$
If $r$ is integrable, the density $\Sigma$ is well-defined and continuous. The CLT for $Z_{\lambda}$ follows from Theorem 1.5 applied to $q=1$, and the lower bound for the variance given by the variance of first chaotic component.
@@
Using the definition of the trace, or equivalently the mixed product property, and Lemma 3.2 applied to the rank one matrix $(1,2i\pi\xi)\otimes(1,-2i\pi\xi)$, one gets thanks to Lemma 4.4 and Lemma 4.5
@@
If $Y$ is not a random wave then this quantity is positive on a set of positive $\nu$-measure, and the conclusion follows from the proof Lemma 3.11. Conversely, if $Y$ is a regular random wave then this quantity cancels for all $\xi$ in the support of $\nu$ and the conclusion of Theorem 1.8 follows from the proof of Lemma 3.12, and in particular from the second point.
@@
It follows from Lemma 4.7 and Lemma 4.8 that
$\operatorname{Tr}\left(f_{2}\Sigma(\xi)f_{2}\Sigma(\xi)\right)\geq\operatorname{Tr}\left(f_{2,1}\Sigma_{1}(\xi)f_{2,1}\Sigma_{1}(\xi)\right)\geq\alpha^{2}\left(-1+\frac{4\pi^{2}}{d}\,\xi^{T}M\xi\right)^{2}|\psi_{1}(\xi)|^{2},$
@@
Without isotropy assumption, 4.10 and Lemma 4.12 implies that
$\operatorname{Tr}\left(f_{2}\tilde{\Sigma}(\xi)f_{2}\tilde{\Sigma}(\xi)\right)=\omega(\xi)^{2}\alpha^{2}\left(-4\pi^{2}\|\xi\|^{2}+S(\xi)M\overline{S(\xi)}\right)^{2}.$
The conclusion follows in a similar fashion as in the previous case. If the field $F$ has an isotropic distribution, then one can explicit this quantity, thanks to Lemma 4.10. One has
$\operatorname{Tr}\left(f_{2}\tilde{\Sigma}(\xi)f_{2}\tilde{\Sigma}(\xi)\right)$ $=\omega(\xi)^{2}f_{2}\left(2i\pi\xi,-\frac{4\pi^{2}}{\sqrt{2\beta}}\left(\xi\otimes\xi-\frac{\gamma}{d}\|\xi\|^{2}\mathrm{Id}\right)\right)^{2}$
$=\omega(\xi)^{2}\left[-4\pi^{2}\|\xi\|^{2}\alpha+\frac{(4\pi^{2})^{2}\|\xi\|^{2}}{2\beta}\left[A\left(1-\frac{2\omega}{2}+\frac{\omega^{2}}{d}\right)+B\left(1-\omega\right)^{2}\right]\right]^{2}$
$=\hat{q}(\xi)^{2}(4\pi^{2}\|\xi\|^{2}\alpha)^{2}\left[-1+\frac{4\pi^{2}\|\xi\|^{2}}{d}\frac{\beta_{0}}{\beta}\right]^{2}$ (21)
The conclusion of Theorem 1.8 follows from the same argument as in the case of hypothesis (H1), using the fact that $\beta=\beta_{0}$ if and only if $F$ is the isotropic random wave model.
@@
For $u$-level with $u\neq 0$, one can prove by similar arguments that there is no second chaos cancellation for the isotropic random waves model. Indeed, Equation (5.2.3) becomes,
@@
$\bullet$ Assume that $Y$ is a symmetric regular random wave:
Let $\mu$ be the spectral measure of the process $(Y,\nabla Y)$ taking values in the space of Hermitian forms on $W=U\otimes(\mathbb{R}\times V)$. Let $\varsigma$ be the density of the spectral measure $\psi$ with respect to the spherical measure $\sigma$. Since the covariance function of $Y$ is symmetric, the spectral measure $\psi$ has no imaginary component and is thus a symmetric measure (w.r.t $\xi$), so is the density $\varsigma$.
We want to apply Theorem 1.4 with $q=2$, and conclude that the variance converges to a limit positive constant, with a logarithmic correction in the case $d=2$. From Lemma 4.4, the spectral measure $\mu$ has a density w.r.t the spherical measure $\sigma$ given by
$\mathrm{d}\mu=\varsigma(\xi)\otimes\left[(1,2i\pi\xi)\otimes(1,-2i\pi\xi)\right]\mathrm{d}\sigma,$
Thanks to Lemma 4.14, the measure $\mu*\mu$ has a density $\Psi$ w.r.t the measure $\sigma*\sigma$, which is bounded and polar-continuous. The variance of the fourth chaos has the expression
$\operatorname{Var}(Z_{\lambda}^{(4)}(\phi))$ $=\frac{1}{24}\int_{V}\left(|\hat{\phi}|^{2}\right)_{\lambda}(\xi)\,\mathrm{d}(\mu*\mu)^{*2}(f_{4})(\xi)$
$=\frac{1}{24}\int_{V}\left(|\hat{\phi}|^{2}\right)_{\lambda}(\xi)\,\mathrm{d}\left[\Psi(\sigma^{*2})*\Psi(\sigma^{*2})\right](f_{4})(\xi)$
@@
The variance asymptotics is explicit thanks to Lemma 4.16. When $d=2$, we have by integration of equivalences,the expression, for positive constants $C,C^{\prime}$
@@
$\lim_{\lambda\rightarrow+\infty}\frac{\operatorname{Var}(Z_{\lambda}^{(4)}(\phi))}{\log(\lambda)}=C\int_{\mathbb{S}^{1}}\operatorname{Tr}(f_{4}\Psi(0,\theta)f_{4}\Psi(0,\theta))\mathrm{d}\theta+C^{\prime}\int_{\|\xi\|=\frac{\sqrt{d}}{\pi}}\operatorname{Tr}(f_{4}\Psi(\xi)f_{4}\Psi(\xi))\mathrm{d}\xi.$
When $d\geq 3$, the limit variance is given by Corollary 1.3 and has expression
$\lim_{\lambda\rightarrow+\infty}\operatorname{Var}(Z_{\lambda}^{(4)}(\phi))=C\int_{V}\operatorname{Tr}(f_{4}\Psi(\xi)f_{4}\Psi(\xi))\left(\frac{\mathrm{d}\sigma*\sigma}{\mathrm{d}\xi}(\xi)\right)^{2}\mathrm{d}\xi.$
In both cases, we can conclude that the limit constant is positive if we are able to show that for $\theta\in\mathbb{S}^{d-1}$ the quantity $\operatorname{Tr}(f_{4}\Psi(0,\theta)f_{4}\Psi(0,\theta))$ is positive. According to Lemma 4.14,
$\Psi(0,\theta)=\int_{\mathbb{S}^{d-1}}\left(\varsigma(\xi)\otimes\left[(1,2i\pi\xi)\otimes(1,-2i\pi\xi)\right]\right)^{\otimes 2}\mathrm{d}\sigma_{\theta}(\xi),$
where $\sigma_{\theta}$ is the uniform probability measure on the $d-2$-sphere on $V$ with normal vector $\theta$. The map $\Psi(0,\theta)$ can be seen as symmetric bilinear map over $S^{2}(W^{*})$, as the fourth chaotic component $f_{4}$, see the paragraph above Lemma 4.6. In particular, it follows from the symmetry of the integrand in Equation (5.2.4) that the component associated to the subspace $U\otimes(U\otimes V)$ of $S^{2}(W^{*})$ defined in Equation (4.2.1) is stable by $\Psi(0,\theta)$. The restriction of $\Psi(0,\theta)$ to this subspace, denoted $\psi^{r}(0,\theta)$, is given by
$\Psi^{r}(0,\theta)=4\pi^{2}\int_{\mathbb{S}^{d-1}}\varsigma(\xi)\otimes\varsigma(\xi)\otimes\left(\xi\otimes\xi\right)\mathrm{d}\sigma_{\theta}(\xi).$
We then have , given the expression of $f_{4}^{r}$ given by Lemma 4.6, and the mixed product property
$\operatorname{Tr}(f_{4}\Psi(0,\theta)f_{4}\Psi(0,\theta))$ $\geq\operatorname{Tr}(f_{4}^{r}\Psi^{r}(0,\theta)f_{4}^{r}\Psi^{r}(0,\theta))$
$\geq\left(\frac{\alpha}{d}\right)^{2}(4\pi^{2})^{2}\int_{\mathbb{S}^{d-1}}\int_{\mathbb{S}^{d-1}}\operatorname{Tr}(\varsigma(\xi)\varsigma(\eta))^{2}\langle\sigma,\eta\rangle^{2}\mathrm{d}\sigma_{\theta}(\xi)\mathrm{d}\sigma_{\theta}(\eta)$
The integrand is continuous, and positive for $\xi=\eta$. We conclude that the whole integral is positive, and so is the limit variance.
To prove the asymptotic normality, note when $d\geq 3$ that $\Omega\in L^{4}$. It follows from Theorem 1.5 that the variance of $Z_{\lambda}^{(4+)}(\phi)$ converges to a finite constant, which is positive thanks to the previous computation , and that $Z_{\lambda}^{(4+)}(\phi)$ satisfies a CLT. Since the variance of $Z_{\lambda}^{(2)}(\phi)$ converges to zero, then $Z_{\lambda}^{(2)}(\phi)$ converges in probability to zero and the CLT for $Z_{\lambda}(\phi)$ follows from Slutsky’s Theorem.
@@
In the case $d=2$, the CLT for $Z_{\lambda}^{(4)}(\phi)$ follows directly from the Fourth Moment theorem, since $\Omega\in L^{6}$ implies that the contractions tends to zero. The variance of $Z_{\lambda}^{(6+)}(\phi)$ converges to a constant by Theorem 1.5, and so does the variance of $Z_{\lambda}(\phi)-Z_{\lambda}^{(4)}(\phi)$. The CLT for $Z_{\lambda}(\phi)$ then follows from Slutsky’s Theorem.
$\bullet$ Assume that $Y$ is the gradient of a real isotropic random wave:
The proof is very similar to the previous case with minor adaptations. Let $\mu$ be the spectral measure of the process $(\nabla F,TF)$ taking values in the space of Hermitian forms on $W=V\otimes(S^{2}(V^{*}))$. From Lemma 4.11, the spectral measure $\mu$ has a density w.r.t the spherical measure $\sigma$ given by
$\mathrm{d}\mu=\left[2i\pi\xi,-\frac{4\pi^{2}}{\sqrt{2\beta}}\left(\xi\otimes\xi-\frac{\gamma}{d}\|\xi\|^{2}\mathrm{Id}\right)\right]\otimes\left[-2i\pi\xi,-\frac{4\pi^{2}}{\sqrt{2\beta}}\left(\xi\otimes\xi-\frac{\gamma}{d}\|\xi\|^{2}\mathrm{Id}\right)\right]\mathrm{d}\sigma.$
Thanks to Lemma 4.14, the measure $\mu*\mu$ has a density $\Psi$ w.r.t the measure $\sigma*\sigma$, which is bounded and polar-continuous. As in the previous paragraph, we can conclude that the limit constant is positive if we are able to show that for $\theta\in S^{d-1}$ the quantity $\operatorname{Tr}(f_{4}\Psi(0,\theta)f_{4}\Psi(0,\theta))$ is positive. According to Lemma 4.14,
$\Psi(0,\theta)=\int\left(\left[2i\pi\xi,-\frac{4\pi^{2}}{\sqrt{2\beta}}\left(\xi\otimes\xi-\frac{\gamma}{d}\|\xi\|^{2}\mathrm{Id}\right)\right]\!\otimes\!\left[-2i\pi\xi,-\frac{4\pi^{2}}{\sqrt{2\beta}}\left(\xi\otimes\xi-\frac{\gamma}{d}\|\xi\|^{2}\mathrm{Id}\right)\right]\right)^{\otimes 2}\!\!\!\!\!\!\!\mathrm{d}\sigma_{\theta}(\xi),$
where $\sigma_{\theta}$ is the uniform probability measure on the $d-2$-sphere on $V$ with normal vector $\theta$. The map $\Psi(0,\theta)$ can be seen as symmetric bilinear map over $S^{2}(W^{*})$, as the fourth chaotic component $f_{4}$, see the paragraph above Lemma 4.13. In particular, it follows from the symmetry of the integrand in Equation (5.2.4) that the component associated to the subspace $V\otimes S^{2}(V)$ of $S^{2}(W^{*})$ defined in Equation (4.2.3) is stable by $\Psi(0,\theta)$. The restriction of $\Psi(0,\theta)$ to this subspace, denoted $\psi^{r}(0,\theta)$, is given by
$\Psi^{r}(0,\theta)=\frac{(4\pi)^{3}}{2\beta}\int_{\mathbb{S}^{d-1}}\left[\xi\otimes\left(\xi\otimes\xi-\frac{\gamma}{4\pi^{2}}\mathrm{Id}\right)\right]^{\otimes 2}\mathrm{d}\sigma_{\theta}(\xi).$
We then have
$\operatorname{Tr}(f_{4}\Psi(0,\theta)f_{4}\Psi(0,\theta))\geq\langle f_{4}^{r}\Psi^{r}(0,\theta)f_{4}^{r}\Psi^{r}(0,\theta))$
$\geq\left(\frac{\alpha}{d}\right)^{2}(4\pi^{2})^{3}\int_{\mathbb{S}^{d-1}}\int_{\mathbb{S}^{d-1}}f_{4}\left(\xi\otimes\left(\xi\otimes\xi-\frac{\gamma}{d}\|\xi\|^{2}\mathrm{Id}\right),\xi\otimes\left(\eta\otimes\eta-\frac{\gamma}{d}\|\xi\|^{2}\mathrm{Id}\right)\right)\mathrm{d}\sigma_{\theta}(\xi)\mathrm{d}\sigma_{\theta}(\eta).$
The integrand is continuous, and when $\xi=\eta$ its value is given, thanks to the expression of $f_{4}^{r}$ given by Lemma 4.13, by
$(\|\xi\|^{2})^{2}\left[A\operatorname{Tr}\left(\left(\xi\otimes\xi-\frac{\gamma}{4\pi^{2}}\mathrm{Id}\right)^{2}\right)+B\operatorname{Tr}\left(\xi\otimes\xi-\frac{\gamma}{4\pi^{2}}\mathrm{Id}\right)^{2}\right]^{2}=\|\xi\|^{4}\left(\frac{2\beta_{0}}{d}\alpha\right)^{2},$
where the last equality follows from the same steps as in Equation (21). We conclude that the whole integral is positive, and so is the limit variance. The asymptotic normality is in all point similar to the previous paragraph.